# Thermodynamic and observational implications of black holes in toroidal geometry

### v2

@@ III.4 PVP-V Diagram and Isotherm compressibility
In the context of Rènyi entropy, we utilize Eq. (18) to compute the thermodynamic pressure in terms of the conjugate temperature, Volume, and charge, which is given as
@@
Figure 2: PVP-V (left panel) and isotherm compressbility $K_{T}$ (right panel) in terms of the Hawking-Bekenstein, exponential corrected and Rényi entropy by utilizing $P=1,~Q=1,~\alpha=1$ (in Rènyi entropy case). We get various curves for different entropy models to notice the effect of these entropies on PVP-V (left panel) and $K_{T}$ (right panel), such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rènyi entropies are presented as red solid, blue dotted, and black dotted-dashed curves, respectively.
@@
PVP-V diagram is considered to be one of the methods to analyze the stability of the BH system. In Fig. 2 left panel, we depict the behavior of PVP-V diagram by utilizing Q=1,P=1Q=1,~P=1 and $\alpha=1$ while we obtain different trajectories in terms of respective entropy models such as the Hawking-Bekenstein (red solid curve), exponential corrected entropy (blue dotted curve) and the Rènyi entropy (black dotted-dashed curve). It observes that pressure $(P)$ smoothly decreases as the volume $(V)$ grows for the Hawking-Bekenstein (red solid curve) and Rènyi entropy (black dotted-dashed curve). This smooth behavior indicates the absence of a possible first-order phase transition, such as the small and large BH transitions corresponding to the Van-dar-Waals fluctuation in the PVP-V diagram. Therefore, this signifies that the BH remains thermodynamically stable, undergoing no abrupt transitions in its internal configuration. However, in the context of exponential corrected entropy (blue dotted curve), we observe that initially, pressure decreases for small values of volume. The pressure begins to increase as the volume increases, which indicates a possible phase transition in the thermodynamic system. In the right panel of Fig. 2, we demonstrate the isotherm compressibility $(K_{T})$ in terms of the Hawking-Bekenstein (red solid curve), exponential corrected entropy (blue dotted curve), and Rènyi entropies (black dotted-dashed curve) by plugging T=1,Q=1T=1,~Q=1 and $\alpha=1$ (Rènyi entropy case). As we mentioned earlier in the PVP-V diagram (left panel), it also validates from the behavior of isotherm compressibility $(K_{T})$ in the form of exponential corrected entropy (blue dotted curve) that there is a possible phase transition while in cases of the Hawking-Bekenstein entropy (red solid curve) and the Rènyi entropy (black dotted-dashed curve) shows increasing behavior for all the values of volume.
@@ III.5 Helmholtz Free Energy
Figure 3: HFE in terms of the Hawking-Bekenstein, exponential corrected, and Rényi entropy by substituting $P=1,~Q=1,~\alpha=1$ (in Rènyi entropy case). We get various curves for different entropy models to observe the effect of these entropies on HFE, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rènyi entropies are depicted as red solid, blue dotted, and black dotted-dashed curves, respectively.
We investigate HFE for the charged torus-like BHs in terms of the Hawking-Bekenstein, exponential corrected, and Rènyi entropies, which is considered one of the methods associated with BHs’ global stability. Behavior of HFE as the function of Hawking-Bekenstein entropy (red solid curve), exponential corrected (blue dotted curve) and Rènyi entropy (black dotted-dashed curve) by setting Q=1,P=1Q=1,~P=1 and $\alpha=1$ is depicted in Fig. 3. It shows that initially HFE decreases with positive behavior, but as entropy increases, the behavior of HFE decreases with negative behavior for both the Hawking-Bekenstein and Rènyi entropies. This fluctuating behavior suggests that charged torus-like BHs are not stable in terms of the Hawking-Bekenstein (red solid curve) and Rènyi (black dotted-dashed curve) entropies. Surprisingly, HFE shows positive behavior for all the ranges of entropy, which suggests stability of HFE for the charged torus-like BHs. Furthermore, this indicates that the thermodynamic behavior of BHs changes based on the entropy formalism used, with both Hawking-Bekenstein and Rényi entropies becoming unstable under specific conditions, such as varying charge or cosmological constant, suggesting limitations in their applicability within particular physical scenarios.
@@ III.6 Gibbs Free Energy
Similarly, by inserting Eqs. (13), (18) in Eq. (33), we calculate GFE in the form of Rènyi entropy, which is given as
@@
Figure 4: GFE in terms of the Hawking-Bekenstein, exponential corrected, and Rényi entropy by utilizing $P=1,~Q=1,~\alpha=1$ (in Rènyi entropy case). We get various curves for different entropy models to notice the impact of these entropies on GFE, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rènyi entropies are demonstrated as red solid, blue dotted, and black dotted-dashed curves, respectively.
In Fig. 4, we present the graphical behavior of GFE as a function of the Hawking-Bekenstein, exponential corrected entropy, and Rènyi entropies by putting Q=1,p=1Q=1,~p=1 and $\alpha=1$. One can notice that initially, GFE in terms of the Rènyi entropy (black dotted-dashed curve) decreases with positive behavior, but as the entropy increases, GFE continuously decreases with a negative trend, which indicates that in the Rènyi entropy case, the behavior of GFE is not stable. However, for both the Hawking-Bekenstein (red solid curve) and the exponential-corrected (blue dotted curve) entropies, GFE initially decreases, but as entropy increases, it begins to rise, indicating its stability across all entropy ranges.
Furthermore, the comparative analysis of all the above-mentioned thermodynamic quantities reveals that the exponential corrected entropy offers a more intricate phase structure, characterized by two specific heat ZPs absent in the Hawking-Bekenstein and Rènyi cases. This distinction reveals a more profound dependence of the microstructure on the horizon’s topology, which we explore geometrically in the following section by employing thermodynamic curvature.
@@ IV Thermodynamic Geometry in terms of Various entropy frameworks
Here, we mention that the Ricci curvature scalar obtained from the Weinhold metric formalism in terms of the Hawking-Bekenstein, exponential corrected and Rènyi entropies are zero (this signifies that the equilibrium manifold is flat in this metric formulation) which allows us to adopt another metric formalism that we employ in this study, is the Ruppeiner metric to investigate the coincidence of the divergence of the Ricci curvature scalar with the ZPs of heat capacity. In this metric formalism, we basically modify the Weinnhold metric given in Eq. (41) by multiplying it with the inverse temperature. The purpose of this scaling, from a physical perspective, is to ensure the consistency of the fluctuation theory underlying the Ruppeiner metric with temperature-dependent thermodynamic probability distributions. For example, the inclusion of inverse temperature enhances our understanding of curvature singularities as indicators of critical behavior or phase transitions in the context of BH thermodynamics. Therefore, the inverse temperature serves not just as a mathematical term but as a thermodynamically significant quantity that influences the structure of fluctuation geometry and reveals key aspects of stability and microscopic behavior. Thereby, one can define the Ruppeiner metric formalism given in Refs. Ruppeiner:2008kd ; Ruppeiner:1995zz ; Akbar:2011qw ; Soroushfar:2020wch , whose explicit form is given as
@@
Similarly, one can easily derive the Ricci curvature of the Ruppeiner formalism in the presence of Rènyi entropy by substituting Eqs. (13) and (18) into Eq. (41), which takes the following shape
@@
Figure 5: Ricci curvature in terms of the Hawking-Bekenstein, exponential corrected, and Rényi entropies by inserting $P=1,~Q=1,~\alpha=1$ (in Rènyi entropy case). We obtain different curves for different entropy models to witness the impact of these entropies on curvature scalar $R^{\mathrm{Rup}}$, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rènyi entropies are illustrated as red solid, blue dotted, and black dotted-dashed curves, respectively.
One can obtain valuable insights regarding the possible phase transition and the microscopic interaction between the particles of the BHs by examining the Ricci curvature scalar. For example, a negative scalar curvature indicates strong attractive microscopic interactions, whereas a positive value suggests repulsive interactions between the particles. When the curvature is flat, it typically denotes either an ideal gas-like non-interacting system or a perfectly balanced interaction regime. The behavior of the curvature scalar obtained from the aforementioned formalism in various entropy frameworks by substituting Q=1,P=1Q=1,~P=1 and $\alpha=1$ is presented in Fig. 5. The trajectories that we plot in Fig. 5 are the Ricci scalar in terms of entropies such as the Hawking-Bekenstein (red solid curve), exponential corrected (blue dotted curve), and Rènyi entropies (black dotted-dashed curve). It observes that the ZPs of heat capacity also coincide with the divergence of the Ricci curvature scalar. In the Hawking-Bekenstein and Rènyi entropies cases, the Ricci curvature scalar $R^{\mathrm{Rup}}$ divergence coincides with ZP of heat capacity at $2.2214$ and $1.1698$, respectively. Similarly, in the case of the exponential corrected entropy, the divergence of $R^{\mathrm{Rup}}$ coincides with both ZPs of heat capacity at $1$ and $3.4674$.
@@ V Sparsity of Hawking radiations and Emission of Energy through Different Entropy Frameworks
(c) Rènyi entropy with $Q$ variations.
(d) Rènyi entropy with $\alpha$ variations.
Figure 6: Sparsity $\eta$ as the function of different entropies such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies by setting $P=1,~\alpha=1$ (in Rèny entropy case) and we substitute $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve) for different curves, in Figs. (6(a))-(6(c)). However, in Fig. (6(d)), we insert Q=1,P=1Q=1,~P=1 and for red solid, blue dotted, and black dotted-dashed curves we utilize $\alpha=0.2,~0.6,~1$, respectively.
Let us mention here that in this analysis of sparsity, we employed the conjugate temperature corresponding to the entropy framework to maintain consistency. To analyze the influence of BH system parameters on Hawking sparsity using different entropy approaches, Fig. 6 presents $\eta$ as a function of entropy $S$ relationship at $P=1$ and different trajectories corresponding to charge $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve), and $1$ (black dotted-dashed curve). We present the sparsity of Hawking radiation by using Hawking-Bekenstein entropy in Fig. 6(6(a)) and exponential corrected entropy in Fig. 6(6(b)). Similarly, we present $\eta$ as the function of entropy by putting Q=0.08,0.5,1Q=0.08,~0.5,~1 for Rènyi entropy in Fig. 6(6(c)) while in Fig. 6(6(d)) we employ different values of non-extensive parameter $\alpha=0.2,~0.6,~1$. It observes that as entropy increases, it leads to a reduction in sparsity, which is not the case for Schwarzschild BHs. Also, for smaller values of entropy, sparsity $\eta$ increases more than the Schwarzschild one, which indicates that, during this phase of evaporation, the BH emits radiation with a lower intensity than Hawking radiation. As $S$ increases, $\eta$ steadily decreases, and approaches zero in the limit. In the case of exponential corrected entropy, $\eta$ decreases as the entropy increases, but after some values of $S$, it shows constant behavior, which is quite evident in Fig. 6(6(b)). Significant changes in the decay behavior arise either due to modifications in the BH parameter space or from adopting different entropy formalisms.
@@
(c) Rènyi entropy with $Q$ variations.
(d) Rènyi entropy with $\alpha$ variations.
Figure 7: The rate of energy emission $\varepsilon_{\omega t}$ as the function of different entropies such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies by setting $P=1,~\omega=1,~\alpha=1$ (in Rèny entropy case) and for different curves we put $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve), in Figs. (7(a))-(7(c)). But for Fig. (7(d)), we put Q=1,P=1Q=1,~P=1 and $\alpha=0.01,~0.5,~1$ for red solid, blue dotted, and black dotted-dashed curves, respectively.
where $T$ is the temperature, which in our analysis corresponds to the different entropy frameworks such as the Hawking-Bekenstein, exponential corrected, and Rènyi entropies. We plot the rate of energy emission $\varepsilon_{\omega t}$ in the form of the Hawking-Bekenstein, exponential corrected, and Rènyi entropies in Fig. 7 by setting $\alpha=1$ and we obtain different trajectories for $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve), and $Q=1$ (black dotted-dashed curve) for Figs. 7(7(a)-7(c)) while in Fig. 7(7(d)), we put $Q=1,~\omega=1,~P=1$ and $\alpha=0.2$ (red solid curve), $\alpha=0.6$ (blue dotted curve) and $\alpha=1$ ( black dotted-dashed dashed curve). For a charged torus-like BH, it is noticed that the rate of energy emission is increasing initially with a negative trend for small ranges of entropy, but as the entropy increases, the values of $\varepsilon_{\omega t}$ also increase with a positive trend. While in the case of exponential corrected entropy, we observe that the rate of emission energy is increasing for small values of entropy with a positive trend, and for large values of entropy, the rate of emission energy slightly decreases, and after that, it becomes constant, which we can clearly observe in Fig. 7(7(b)). However, the behavior of $\varepsilon_{\omega t}$ in Figs. 7(7(a),7(c),7(d)) demonstrates that for large values of entropy, BH radiates permanently. The variation in the $\varepsilon_{\omega t}$ led BHs in high-energy configurations to emit radiation of low frequency with large values of entropy, which makes their identification very complicated. These phenomena are quite significant in the observational astronomy of BH evaporation for differentiating between the traditional general relativity and other gravity frameworks.

### v3

@@ III.4 PVP-V Diagram and Isotherm compressibility
In the context of Rényi entropy, we utilize Eq. (18) to compute the thermodynamic pressure in terms of the conjugate temperature, Volume, and charge, which is given as
@@
Figure 2: PVP-V (left panel) and isotherm compressbility $K_{T}$ (right panel) in terms of the Hawking-Bekenstein, exponential corrected and Rényi entropy by utilizing $P=1,~Q=1,~\alpha=1$ (in Rényi entropy case). We get various curves for different entropy models to notice the effect of these entropies on PVP-V (left panel) and $K_{T}$ (right panel), such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rényi entropies are presented as red solid, blue dotted, and black dotted-dashed curves, respectively.
@@
PVP-V diagram is considered to be one of the methods to analyze the stability of the BH system. In Fig. 2 left panel, we depict the behavior of PVP-V diagram by utilizing Q=1,P=1Q=1,~P=1 and $\alpha=1$ while we obtain different trajectories in terms of respective entropy models such as the Hawking-Bekenstein (red solid curve), exponential corrected entropy (blue dotted curve) and the Rényi entropy (black dotted-dashed curve). It observes that pressure $(P)$ smoothly decreases as the volume $(V)$ grows for the Hawking-Bekenstein (red solid curve) and Rényi entropy (black dotted-dashed curve). This smooth behavior indicates the absence of a possible first-order phase transition, such as the small and large BH transitions corresponding to the Van-dar-Waals fluctuation in the PVP-V diagram. Therefore, this signifies that the BH remains thermodynamically stable, undergoing no abrupt transitions in its internal configuration. However, in the context of exponential corrected entropy (blue dotted curve), we observe that initially, pressure decreases for small values of volume. The pressure begins to increase as the volume increases, which indicates a possible phase transition in the thermodynamic system. In the right panel of Fig. 2, we demonstrate the isotherm compressibility $(K_{T})$ in terms of the Hawking-Bekenstein (red solid curve), exponential corrected entropy (blue dotted curve), and Rényi entropies (black dotted-dashed curve) by plugging T=1,Q=1T=1,~Q=1 and $\alpha=1$ (Rényi entropy case). As we mentioned earlier in the PVP-V diagram (left panel), it also validates from the behavior of isotherm compressibility $(K_{T})$ in the form of exponential corrected entropy (blue dotted curve) that there is a possible phase transition while in cases of the Hawking-Bekenstein entropy (red solid curve) and the Rényi entropy (black dotted-dashed curve) shows increasing behavior for all the values of volume.
@@ III.5 Helmholtz Free Energy
Figure 3: HFE in terms of the Hawking-Bekenstein, exponential corrected, and Rényi entropy by substituting $P=1,~Q=1,~\alpha=1$ (in Rényi entropy case). We get various curves for different entropy models to observe the effect of these entropies on HFE, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rényi entropies are depicted as red solid, blue dotted, and black dotted-dashed curves, respectively.
We investigate HFE for the charged torus-like BHs in terms of the Hawking-Bekenstein, exponential corrected, and Rényi entropies, which is considered one of the methods associated with BHs’ global stability. Behavior of HFE as the function of Hawking-Bekenstein entropy (red solid curve), exponential corrected (blue dotted curve) and Rényi entropy (black dotted-dashed curve) by setting Q=1,P=1Q=1,~P=1 and $\alpha=1$ is depicted in Fig. 3. It shows that initially HFE decreases with positive behavior, but as entropy increases, the behavior of HFE decreases with negative behavior for both the Hawking-Bekenstein and Rényi entropies. This fluctuating behavior suggests that charged torus-like BHs are not stable in terms of the Hawking-Bekenstein (red solid curve) and Rényi (black dotted-dashed curve) entropies. Surprisingly, HFE shows positive behavior for all the ranges of entropy, which suggests stability of HFE for the charged torus-like BHs. Furthermore, this indicates that the thermodynamic behavior of BHs changes based on the entropy formalism used, with both Hawking-Bekenstein and Rényi entropies becoming unstable under specific conditions, such as varying charge or cosmological constant, suggesting limitations in their applicability within particular physical scenarios.
@@ III.6 Gibbs Free Energy
Similarly, by inserting Eqs. (13), (18) in Eq. (33), we calculate GFE in the form of Rényi entropy, which is given as
@@
Figure 4: GFE in terms of the Hawking-Bekenstein, exponential corrected, and Rényi entropy by utilizing $P=1,~Q=1,~\alpha=1$ (in Rényi entropy case). We get various curves for different entropy models to notice the impact of these entropies on GFE, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rényi entropies are demonstrated as red solid, blue dotted, and black dotted-dashed curves, respectively.
In Fig. 4, we present the graphical behavior of GFE as a function of the Hawking-Bekenstein, exponential corrected entropy, and Rényi entropies by putting Q=1,p=1Q=1,~p=1 and $\alpha=1$. One can notice that initially, GFE in terms of the Rényi entropy (black dotted-dashed curve) decreases with positive behavior, but as the entropy increases, GFE continuously decreases with a negative trend, which indicates that in the Rényi entropy case, the behavior of GFE is not stable. However, for both the Hawking-Bekenstein (red solid curve) and the exponential-corrected (blue dotted curve) entropies, GFE initially decreases, but as entropy increases, it begins to rise, indicating its stability across all entropy ranges.
Furthermore, the comparative analysis of all the above-mentioned thermodynamic quantities reveals that the exponential corrected entropy offers a more intricate phase structure, characterized by two specific heat ZPs absent in the Hawking-Bekenstein and Rényi cases. This distinction reveals a more profound dependence of the microstructure on the horizon’s topology, which we explore geometrically in the following section by employing thermodynamic curvature.
@@ IV Thermodynamic Geometry in terms of Various entropy frameworks
Here, we mention that the Ricci curvature scalar obtained from the Weinhold metric formalism in terms of the Hawking-Bekenstein, exponential corrected and Rényi entropies are zero (this signifies that the equilibrium manifold is flat in this metric formulation) which allows us to adopt another metric formalism that we employ in this study, is the Ruppeiner metric to investigate the coincidence of the divergence of the Ricci curvature scalar with the ZPs of heat capacity. In this metric formalism, we basically modify the Weinnhold metric given in Eq. (41) by multiplying it with the inverse temperature. The purpose of this scaling, from a physical perspective, is to ensure the consistency of the fluctuation theory underlying the Ruppeiner metric with temperature-dependent thermodynamic probability distributions. For example, the inclusion of inverse temperature enhances our understanding of curvature singularities as indicators of critical behavior or phase transitions in the context of BH thermodynamics. Therefore, the inverse temperature serves not just as a mathematical term but as a thermodynamically significant quantity that influences the structure of fluctuation geometry and reveals key aspects of stability and microscopic behavior. Thereby, one can define the Ruppeiner metric formalism given in Refs. Ruppeiner:2008kd ; Ruppeiner:1995zz ; Akbar:2011qw ; Soroushfar:2020wch , whose explicit form is given as
@@
Similarly, one can easily derive the Ricci curvature of the Ruppeiner formalism in the presence of Rényi entropy by substituting Eqs. (13) and (18) into Eq. (41), which takes the following shape
@@
Figure 5: Ricci curvature in terms of the Hawking-Bekenstein, exponential corrected, and Rényi entropies by inserting $P=1,~Q=1,~\alpha=1$ (in Rényi entropy case). We obtain different curves for different entropy models to witness the impact of these entropies on curvature scalar $R^{\mathrm{Rup}}$, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rényi entropies are illustrated as red solid, blue dotted, and black dotted-dashed curves, respectively.
One can obtain valuable insights regarding the possible phase transition and the microscopic interaction between the particles of the BHs by examining the Ricci curvature scalar. For example, a negative scalar curvature indicates strong attractive microscopic interactions, whereas a positive value suggests repulsive interactions between the particles. When the curvature is flat, it typically denotes either an ideal gas-like non-interacting system or a perfectly balanced interaction regime. The behavior of the curvature scalar obtained from the aforementioned formalism in various entropy frameworks by substituting Q=1,P=1Q=1,~P=1 and $\alpha=1$ is presented in Fig. 5. The trajectories that we plot in Fig. 5 are the Ricci scalar in terms of entropies such as the Hawking-Bekenstein (red solid curve), exponential corrected (blue dotted curve), and Rényi entropies (black dotted-dashed curve). It observes that the ZPs of heat capacity also coincide with the divergence of the Ricci curvature scalar. In the Hawking-Bekenstein and Rényi entropies cases, the Ricci curvature scalar $R^{\mathrm{Rup}}$ divergence coincides with ZP of heat capacity at $2.2214$ and $1.1698$, respectively. Similarly, in the case of the exponential corrected entropy, the divergence of $R^{\mathrm{Rup}}$ coincides with both ZPs of heat capacity at $1$ and $3.4674$.
@@
IV.1 A detailed thermodynamic comparison of various entropy models
In this subsection, we compare the thermodynamic behavior obtained from the different entropy models along with the classical BH solutions, such as Schwarzschild and RN-AdS BHs. For completeness, we employ the generalized three-parameter entropy (GTPE) model and compare it with the entropy models we discussed in this study. Firstly, we discuss the mathematical expression of GTPE, which is obtained from the generalised entropy frameworks given in Refs. Elizalde:2025iku ; Nojiri:2022aof ; Nojiri:2024zdu , and its expression can be written as
$\mathcal{S}_{3}[\alpha,~\beta,~\gamma]=\frac{1}{\gamma}\left\{\left(1+\frac{S_{\mathrm{HB}}\alpha}{\beta}\right)^{\beta}-1\right\}\,$, (45)
where $\alpha,~\beta,~\gamma\geq 0$ are the free parameters. We mention here that this entropy is the reduced form of generalized entropy given in Refs. Nojiri:2024zdu , which contains six free parameters. Now, if one can choose $\gamma=\alpha$, one can obtain the Sharma-Mittal entropy jahromi2018generalized ; Nojiri:2024zdu , while if one chooses $\alpha\to\infty$, then Eq. (45) leads to the Tsallis and Barrow entropy tsallis1988possible ; Barrow:2020tzx . Similarly, by choosing $\alpha,~\beta\to\infty$ with $\frac{\alpha}{\beta}$ finite, which eventually yields the Rényi entropy Odintsov:2022qnn , and also if we put $\gamma=\alpha,~\beta\to\infty$, then we can obtain the Loop Quantum gravity entropy Majhi:2017zao . Additionally, one can easily recover the Hawking-Bekenstein entropy Bekenstein:1973ur ; Nojiri:2024zdu , by replacing $\alpha_{+}\to\infty,~\beta=1,~\gamma=(\alpha/\beta)^{\beta}$ in Eq. (45).
Table 1: This table presents the thermodynamic quantities by employing the Hawking-Bekenstein (HB), Rényi, and exponential corrected (EC) entropies for the charged torus-like BH, and also its comparison with the thermodynamic quantities of the RN-Ads and Schwarzschild BHs. For a comprehensive analysis, we compared our entropy models with the generalized entropy by incorporating GTPE from Refs. Elizalde:2025iku ; Nojiri:2022aof . Here, we present the thermodynamic quantities like temperature, heat capacity, isothermal compressibility, HFE, GFE, and Ruppeiner Ricci scalar by $C,~K_{T}~\mathcal{F},~\mathcal{G}$, and $R^{\mathrm{Rup}}$, respectively.
BHs CTL BH RN-AdS Schwarzschild
Quantitites HBE Rényi ECE GTPE HBE HBE
$C$ UnStable Unstable UnStable UnStable Unstable Unstable
ZPs 1 1 2 1 1 0
PVP-V Monotonic Monotonic Non-monotonic Monotonic Non-monotonic not defined
$K_{T}$ Positive Positive Divergence Positive Divergence not defined
$\mathcal{F}$ UnStable Unstable Stable Unstable Unstable Unstable
$\mathcal{G}$ Stable Unstable Stable Unstable Stable Stable
$R^{\mathrm{Rup}}$ Coincide Coincide Coincide both ZPs Coincide Coincide No-coincidence
Table 1 summarizes the thermodynamic quantities for the charged torus-like BHs derived employing the Hawking-Bekenstein, Rényi, exponential corrected, and the GTPE model, together with those for RN-AdS and Schwarzschild BHs. It becomes from our analysis that the thermodynamic behavior of the Hawking-Bekenstein, Rényi, and GTPE models is qualitatively similar. In contrast, the exponentially corrected entropy yields a distinct trend in all thermodynamic quantities. For example, exponential corrected entropy leads to two heat capacity ZPs, while the remaining entropy formulations give only one ZP, a behavior further confirmed by the associated Ruppeiner curvature divergences. This comparison indicates that the exponential corrected entropy responds more sensitively to BH microstructure and horizon topology.
@@ V Sparsity of Hawking radiations and Emission of Energy through Different Entropy Frameworks
(c) Rényi entropy with $Q$ variations.
(d) Rényi entropy with $\alpha$ variations.
Figure 6: Sparsity $\eta$ as the function of different entropies such as the Hawking-Bekenstein, exponential corrected, and Rényi entropies by setting $P=1,~\alpha=1$ (in Rény entropy case) and we substitute $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve) for different curves, in Figs. (6(a))-(6(c)). However, in Fig. (6(d)), we insert Q=1,P=1Q=1,~P=1 and for red solid, blue dotted, and black dotted-dashed curves we utilize $\alpha=0.2,~0.6,~1$, respectively.
Let us mention here that in this analysis of sparsity, we employed the conjugate temperature corresponding to the entropy framework to maintain consistency. To analyze the influence of BH system parameters on Hawking sparsity using different entropy approaches, Fig. 6 presents $\eta$ as a function of entropy $S$ relationship at $P=1$ and different trajectories corresponding to charge $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve), and $1$ (black dotted-dashed curve). We present the sparsity of Hawking radiation by using Hawking-Bekenstein entropy in Fig. 6(6(a)) and exponential corrected entropy in Fig. 6(6(b)). Similarly, we present $\eta$ as the function of entropy by putting Q=0.08,0.5,1Q=0.08,~0.5,~1 for Rényi entropy in Fig. 6(6(c)) while in Fig. 6(6(d)) we employ different values of non-extensive parameter $\alpha=0.2,~0.6,~1$. It observes that as entropy increases, it leads to a reduction in sparsity, which is not the case for Schwarzschild BHs. Also, for smaller values of entropy, sparsity $\eta$ increases more than the Schwarzschild one, which indicates that, during this phase of evaporation, the BH emits radiation with a lower intensity than Hawking radiation. As $S$ increases, $\eta$ steadily decreases, and approaches zero in the limit. In the case of exponential corrected entropy, $\eta$ decreases as the entropy increases, but after some values of $S$, it shows constant behavior, which is quite evident in Fig. 6(6(b)). Significant changes in the decay behavior arise either due to modifications in the BH parameter space or from adopting different entropy formalisms.
@@
(c) Rényi entropy with $Q$ variations.
(d) Rényi entropy with $\alpha$ variations.
Figure 7: The rate of energy emission $\varepsilon_{\omega t}$ as the function of different entropies such as the Hawking-Bekenstein, exponential corrected, and Rényi entropies by setting $P=1,~\omega=1,~\alpha=1$ (in Rèny entropy case) and for different curves we put $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve) and $Q=1$ (black dotted-dashed curve), in Figs. (7(a))-(7(c)). But for Fig. (7(d)), we put Q=1,P=1Q=1,~P=1 and $\alpha=0.01,~0.5,~1$ for red solid, blue dotted, and black dotted-dashed curves, respectively.
where $T$ is the temperature, which in our analysis corresponds to the different entropy frameworks such as the Hawking-Bekenstein, exponential corrected, and Rényi entropies. We plot the rate of energy emission $\varepsilon_{\omega t}$ in the form of the Hawking-Bekenstein, exponential corrected, and Rényi entropies in Fig. 7 by setting $\alpha=1$ and we obtain different trajectories for $Q=0.08$ (red solid curve), $Q=0.5$ (blue dotted curve), and $Q=1$ (black dotted-dashed curve) for Figs. 7(7(a)-7(c)) while in Fig. 7(7(d)), we put $Q=1,~\omega=1,~P=1$ and $\alpha=0.2$ (red solid curve), $\alpha=0.6$ (blue dotted curve) and $\alpha=1$ ( black dotted-dashed dashed curve). For a charged torus-like BH, it is noticed that the rate of energy emission is increasing initially with a negative trend for small ranges of entropy, but as the entropy increases, the values of $\varepsilon_{\omega t}$ also increase with a positive trend. While in the case of exponential corrected entropy, we observe that the rate of emission energy is increasing for small values of entropy with a positive trend, and for large values of entropy, the rate of emission energy slightly decreases, and after that, it becomes constant, which we can clearly observe in Fig. 7(7(b)). However, the behavior of $\varepsilon_{\omega t}$ in Figs. 7(7(a),7(c),7(d)) demonstrates that for large values of entropy, BH radiates permanently. The variation in the $\varepsilon_{\omega t}$ led BHs in high-energy configurations to emit radiation of low frequency with large values of entropy, which makes their identification very complicated. These phenomena are quite significant in the observational astronomy of BH evaporation for differentiating between the traditional general relativity and other gravity frameworks.